### v1

# Emotional Strain and Frustration Drivers in LLM-Human Interactions in Software Engineering

Abstract
Large Language Models (LLMs) are increasingly integrated into various daily tasks in Software Engineering. Even though LLMs are widely used, some interactions can lead to unexpected challenges, which can cause emotions that can evolve into frustration. Frustration can negatively impact the engineer’s productivity and well-being if it escalates into stress and burnout, particularly given the demanding nature of Software Engineering. We explored software engineers’ emotional responses, specifically strains, and identified common causes of frustration when interacting with LLMs at work. We collected 62 survey responses from software practitioners and academics across various software companies and universities, using open-ended, closed-ended, and Likert-scale questions. Then, we analysed the survey responses using content analysis and descriptive statistics.
Our findings revealed that a majority of software engineers experience frustrations or other emotions that can evolve into frustrations. More specifically, we found that the leading causes of frustrations aligned with the engineers’ expectations and were related mainly to the responses’ correctness and reliability. Moreover, we noticed that another major frustration trigger comes from uncritical problems, such as the LLM failing to meet personal preferences (e.g., the output structure) since it would require the engineer to refactor or make more modifications to the LLM output before using it.
This study not only explores frustration triggers when using LLMs but also proposes guidelines to improve the software engineers’ experience, aiming to minimise long-term consequences on mental health.
Index Terms:
Software Engineering, Large Language Models (LLMs), Frustration, Emotions
I Introduction
Software Engineering is full of challenges, from fixing bugs to dealing with changing requirements. Recently, Large Language Models (LLMs) like ChatGPT and GitHub Copilot have been used by software engineers to assist them in performing various tasks including code generation, and quality assurance [asare2023github, lubos2024leveraging]. Current research focuses on understanding how practitioners aim to increase their productivity and make their work process more efficient by targeting LLMs to automate the generation of software artifacts or receive guidance on how to solve certain problems [khojah2024beyond, storey2016disrupting].
However, when certain challenges and limitations of LLMs hinder their effectiveness (e.g., not arriving at a helpful response), engineers may experience frustration [weisz2022better]. Frustrations cause by LLMs as technologies are a component of Techno-stress that impact the engineers’ overall workflow, well-being, and productivity [tarafdar2007impact, muller2023can].
While frustrations have gotten little attention in LLM4SE research, we argue that understanding the causes of frustrations when using LLMs for software-related tasks is the first step to minimizing them and thus improving the productivity and well-being of practitioners both in the software engineering industry and academia.
In this study, we conduct an exploratory study to empirically investigate the causes of frustrations when software engineers interact with LLMs and propose ways to work towards improving productivity and well-being. We focus on the following research questions:
RQ1: What are the triggers or sources of frustration of software engineers when using LLMs?
Our study presents four main categories that can cause frustrating emotions for the software engineer. We found that the main cause of frustration is when the software engineers receive an unhelpful or incorrect answer, followed by misunderstanding the intention, failing to meet personal preferences and other limitations of LLMs. We argue that these categories are specific to software engineering since it is a domain that requires precise, context-aware, and technically accurate responses that can be directly applied, like using LLMs for code generation.
RQ2: How can the frustrating experience impact motivation?
We report that motivation can be temporarily reduced when certain expectations are unmet and frustration is caused. However, interestingly, we found that the software engineers’ motivation to continue with their tasks mostly remains unaffected by a frustrating interaction with LLMs. This was due to the urgency to complete the task (e.g., short sprints and deadlines), and the availability of alternatives (e.g., search engines or personal expertise).
RQ3: How can the experience be less frustrating?
We provide recommendations on areas to improve in LLMs that can be considered by chatbot designers. These recommendations align with the categories of different expectations of software engineers from LLMs. In addition, we argue for the need to provide training and raise awareness among software engineers (in academia and industry) in order to manage and minimize frustrations.
II Background and Related Work
This sections presents the conception framework of this study, as well as, previous work done on the topic and the the research gap we are trying to address.
II-1 Large language models for Software Engineering
With the rise of Large Language Models (LLMs), many industries have been undergoing a transition towards using them for various activities. In Software Engineering, recent research has shown that LLMs have the potential to support practitioners in a variety of tasks including, but not limited to implementation, testing, requirements engineering [dakhel2023github, wang2024software, norheim2024challenges, khojah2024beyond]. Moreover, despite the challenges LLMs impose on academia (e.g., bias and hallucination), they provide many opportunities for researchers and educators as assistants with creating study guides and academic writing [meyer2023chatgpt].
Despite this extensive use and the time that LLMs can save when automating tasks, integrating LLMs into SE practices can lead to user frustration. Researchers have noted that while LLMs assist software practitioners, they may sometimes generate errors, causing them to spend additional time solving errors or seeking clarification. This can also trigger negative emotions, such as frustration [weisz2022better].
II-A Emotions
In this study, we considered the definition by the American Psychological Association [apa_dictionary_emotion], “a complex reaction pattern, involving experiential, behavioural, and physiological elements, by which an individual attempts to deal with a personally significant matter or event”. We acknowledge the diversity of the conceptual definitions of emotions, we selected this definition for its scope, which integrates experiential, behavioural, and physiological dimensions. The Feeling Wheel, developed by Gloria Willcox [willcox1982feeling], is a psychological tool designed to facilitate identifying and expressing emotions. It is widely used in therapeutic settings and personal development to help individuals articulate their emotional states more precisely. The tool organises emotions into a concentric structure, with six core emotions—happy, sad, angry, scared, strong, and calm—situated at the center of the wheel, see the wheel in Figure1. As one moves outward from the core, these primary emotions subdivide into more specific and nuanced feelings, allowing for a more granular understanding of emotional experiences.
Figure 1: Willcox’s Emotions Wheel [neff2024feelings]
In research contexts, the Feeling Wheel has been employed to categorise participant-reported emotions, enabling a systematic approach to the analysis of emotional expression. In this study, we used the adaptation done by [neff_emotional_2023] to explore how individuals experience their interaction with LLMs.
II-A1 Emotional Responses when Using Technology
Salanova [salanova2014technostress] defines Technostrain as a negative psychological experience characterised by (1) heightened anxiety and fatigue (affective dimension), (2) scepticism (attitudinal dimension), and (3) a sense of inefficacy (cognitive dimension) associated with technology use. Aligned with this, Muller et al. [muller2023can] researched techno-frustration, a component of technostrain, which refers to the psychological strain caused by the disorganised or inefficient use of ICT. Techno-frustration describes the experience of feeling discouraged, uncertain, stressed, confused, and upset as a result of using ICT. Such psychological strain can lead to decreased job satisfaction or an increased risk of burnout [srivastava2015technostress]. Muller et al.’s work is the closest to our study; however, we focus on a specific interaction, LLM-user interaction, which is still underexplored despite the rapid integration of LLMs in software engineering.
Regarding research on how technology impacts users’ emotions, studies have focused on the psychological effects of prolonged technology use, recognising that digital tools influence productivity and mental and emotional states (e.g.,[calvo2014positive, mark2015focused]). This focus reflects an effort to explore the balance between the benefits of technology and the potential strain it places on users, particularly in high-demand environments like software engineering. Examples of those efforts are in the form of interventions (e.g.[montes2024qualifying], [penzenstadler2022take]) or looking for the causes (e.g. [wong2023mental]).
Furthermore, users’ expectations when using technology, particularly LLMs, are crucial to their frustration. Studies have shown that prior experiences with technology can shape users’ expectations, influencing their perceptions of performance and trust [ferreri2023identifying]. External factors, such as marketing and communication from technology developers, also play a critical role in setting these expectations [roscoe2017presentation]. When technology does not meet users’ expectations—whether based on past interactions or external portrayals—frustration can escalate, hindering effective task completion and satisfaction  [srivastava2015technostress]. This study aims to explore whether these established findings apply also to LLM interactions, and how this impacts software engineering tasks, a question currently underexplored in existing literature.
This study addresses a gap in the literature by exploring software engineers’ emotional responses, particularly their frustration when using large language models (LLMs), motivated by the growing use of LLMs in professional tasks.
This research aims to provide insights that can inform improvements in LLM design and user experience, ultimately enhancing software engineering productivity and reducing technostress.
III Methodology
The main goal of our study is to understand software engineers’ frustrating experiences when interacting with LLMs in order to be able to point out potential causes of frustrations and possible mitigation strategies. We are also interested in looking into other emotions than frustration that software engineers feel in different interactions and scenarios. We illustrate the different components of our methodology in Figure 2.
Figure 2: Exploratory Study Process.
III-A Study Design
This study employed an exploratory design to investigate the frustrations experienced by software engineers when using large language models (LLMs). Given the limited existing research on this specific topic, an exploratory approach was deemed appropriate to gain initial insights and identify patterns in how engineers interact with LLMs [stebbins2001exploratory]. Exploratory research is particularly useful when studying relatively new phenomena, as it allows for collecting qualitative and quantitative data without predefined hypotheses, facilitating a broad understanding of the issues at hand. We collected open and closed questions to ensure a comprehensive data set that addresses both specific aspects of LLM usage and participants’ broader, more subjective experiences.
III-B Data Collection
In this study, we used a survey as our main data collection method. We initially designed the survey questions based on previous research about the usage of LLMs in Software Engineering and how it impacts their experience and productivity [khojah2024beyond] as well as general limitations of LLMs [hadi2023survey].
After a pilot study with two PhD students (academia) and two software engineers (industry), we added more questions and refined the survey (available in our Zenodo package [zenodo]). In general, the survey includes (i) demographic questions and questions related to the participants’ experience and familiarity with LLMs, (ii) questions about their usage of LLMs during their work and their expectations from the LLMs, (iii) questions about the participants’ emotions when certain expectations are not met, (iv) open-ended questions about frustrations the participants experienced when using LLMs, and (v) Likert-scales of the level of importance of different LLMs abilities and aspects to minimize frustrating experiences.
We created the survey on Google Forms and distributed it across social platforms (LinkedIn and Facebook). We also sent personal invitations to employees who work across 10 software organizations of different sizes (Startups and large companies) and domains (e.g., automotive and eLearning). We also invited researchers and academics in software engineering conferences (RE’24 and FSE’24) to allow our sample of software engineers to be diverse in terms of countries and domains.
III-C Data Analysis
To analyse the open questions, we used content analysis with an inductive approach following the steps by Erlingsson and Brysiewicz [erlingsson2017hands]. The authors analysed the data systematically, aiming to generate patterns, codes and themes [lune2017qualitative]. We discussed each answer for each question, reaching a data agreement. The codes were then categorised based on similarities and differences [erlingsson2017hands]. For question eight, we used the emotion classification and feeling wheel by Willcox [willcox1982feeling] to categorise and identify the range of emotions expressed by participants, enabling a deeper understanding of their emotional responses during interactions with LLMs. For the Likert and close questions, we use descriptive statistics and create visualisations to improve readability. We could not perform a correlation analysis due to the size of our sample.
IV Results
IV-A Respondents Demographics
The responses in our survey showed a diverse sample of software engineers in terms of ages (See Figure 3), and roles (See Table I). The organizations where across 7 countries distributed over 3 continents and are within the domains: Airline and Aircrafts, Automotive, Game design, Infotainment, eLearning, Cybersecurity, Telecommunication, Trade, and Software Engineering research and education.
Figure 3: The distribution of the ages of the participants.
TABLE I: Participants’ areas and roles.
Area Roles # Participants Total
PhD Student 15
Academia Researcher 7 27
Professor 5
Software Developer 10
Software Engineer 6
Manager 5
AI Engineer 4
Industry Researcher 3 35
Tech Lead 2
Software Designer 2
Software Tester 1
Application Specialist 1
Applied Scientist 1
Regarding the nature of their familiarity with LLMs, the vast majority of our participants (58 out of 62) describe themselves as “familiar” or “very familiar” with LLMs. Our participants also describe the wide range of LLMs that they use at work in Figure 4 with ChatGPT being the most popular one.
IV-B Usage of LLMs in Software Engineering industry and academia
The results show that 56 participants (94.9%) use LLMs occasionally, out of which 38 participants (66%) use them on a weekly or daily basis. Only 3 (5.1%) participants indicated that they rarely use LLMs at work. Regarding the LLMs that the participants use at work, the results (depicted in Figure 4) show that ChatGPT was by far the most used one. Table II shows the tasks for which academia and industry respondents apply them. We considered more than once answer per participant.
Figure 4: LLMs used by our participants.
In industry, respondents use LLMs for programming tasks like code generation, debugging, and optimisation. They also employ LLMs for written communication tasks, such as drafting emails and brainstorming ideas, and for generating and improving text. Additionally, LLMs help users onlearning new technologies and research by providing starting points, best practices, and summaries of lengthy information. Lastly, respondents view LLMs as digital assistants for task management and problem-solving, streamlining workflows and enhancing productivity.
TABLE II: LLM Usage Across Industrial and Academic Tasks
Usage Group Area Total Count
Writing Tasks Academia 13
Programming Tasks Academia 8
Educational Tasks Academia 8
Research Related tasks Academia 5
Non-Critical Tasks Academia 5
Programming Tasks Industry 18
Written Communication Tasks Industry 10
Learning New Concepts Industry 8
Task Management Industry 5
In academia, LLMs are primarily used for writing tasks, including generating drafts, checking grammar, and providing content suggestions. Users find them helpful for managing busy work, such as email writing and idea generation, and for creating initial drafts for refinement. Additionally, participants view LLMs as educational tools, using them to understand new technologies and programming concepts or to assist in teaching. For programming tasks, LLMs help write simple code, debug, and learn new coding concepts, offering initial code snippets and quick insights into technologies.
LLMs are also employed for research-related tasks such as summarising academic papers, generating ideas, and finding references. They assist in translating data, cleaning datasets, and extracting key information from research.
IV-C Emotions during LLM interaction
We analyse the emotional responses described by our participants when interacting with LLMs during work, more specifically when the interaction takes a different path than expected and provide an incorrect response. Since emotional responses are complex and layered rather than straightforward, there was a one-to-many relationship between participants and emotions, where participants described a combination of multiple emotions when describing their experiences.
We used the emotions wheel by Wilcox (See Figure 1) to map participants’ answers to various emotional states, allowing us to categorise and analyse the range of emotions experienced when using LLMs. According to Willcox’s emotions wheel, frustration comes initially from anger, which can develop into annoyance, then eventually frustration. These three emotions make up more than half of the emotions observed among our participants in Figure 5. More specifically, the most common feelings, that is, frustration (30.4%) and annoyance (17.4%), can challenge software engineers’ well-being, and a continuous exposure to these emotions can also create a barrier to smooth workflows and a cumulative risk of emotional strain over time.
Other participants reported feeling indifferent (13%), while some showed disappointment (10.1%) which are related to how the engineers’ expectations are met by the LLM.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 29: “Knowing how LLM[s] work, I typically have lower expectations. So I [don’t] feel as frustrated or disappointed, particularly if I know that the task I asked is not trivial. I just try again in a different way.”
Such expectations come from building knowledge about the LLM being used as well as understanding its capabilities and limitations based on previous interactions.
Less frequent reactions included positive emotions like calmness (5.8%), playfulness (2.9%) and curiosity (1.4%), as well as negative emotions such as anger (2.9%), confusion (2.9%), stress (1.4%), and sadness (1.4%). Some participants also reported feeling thoughtful (1.4%) or even guilty (1.4%) showing the varied and sometimes unexpected emotional responses that emerged. These emotions suggest that interacting with LLMs is not just a functional exchange but an exploratory experience for some software engineers. Curiosity, for example, has attached an investigative mindset often aligned with a trial-and-error approach. Playfulness, meanwhile, shows a willingness to engage with the LLM on a more relaxed, open-ended basis. Guilt, though rare, introduces a new angle, suggesting that some engineers may feel a sense of responsibility or even self-blame if the interaction does not go as expected.
Figure 5: Emotional responses when receiving an incorrect answer. The colors map to Willcox’s emotions wheel in Figure 1.
IV-D Expectations when interacting with LLMs
As shown in Table III, participants’ expectations when using LLMs go beyond mere functionality and extend into deeper aspects of quality, reliability, and versatility. For quality and performance, software engineers expect LLMs to consistently deliver correct and unhallucinated information without error or delay, as inconsistency can erode trust. The participants also indicated that LLMs must be user-friendly, that is, ensure a smooth and easy interaction with the user while ensuring data protection, reflecting participants’ concerns over privacy in sensitive tasks. In terms of understanding, engineers value LLMs that grasp context and intent, highlighting the need for deeper comprehension of nuanced queries. Some added that they expect the LLM to understand the context without the need for an excessive description of the context in the initial prompt, but also ask clarification questions if needed to better understand the intent and the context.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 38: “I expect LLMs to recognize and alert me if they are lacking certain information from my prompt”
When it comes to answer structure, most of the participants who use LLMs for research or formatting text emphasised the importance of clear, organised, and grammatically correct responses. In addition, engineers preferred concise or elaborative answers depending on their needs and the task they want support with. For code-related tasks, engineers preferred that the response starts with the snippet of code that is needed (e.g., to repair a program) followed by an explanation that the engineer can optionally go through. While elaborative answers where perceived beneficial for engineers who used LLMs to learn new concepts or explain certain artifacts (e.g., error message). Information integrity was also expected from LLMs. More specifically, the participants mentioned how the LLM should be able to provide the sources of information that were used to generate the output, as well as the confidence estimation of the accuracy of the output. The participants motivated these two transparency criteria as a means to enable the verification of the information and support decision making (e.g., deciding how to utilize the output). Finally, in terms of versatility, our findings showed that integration with other tools and adaptability to different workflows are increasingly expected by engineers in industry as the adoption of LLMs is increasing in software organisations. This adaptability allows the LLMs to fit into varied contexts and communication preferences, signalling a strong need for flexibility.
TABLE III: Users’ Expectations when Using LLMs
Themes Categories
Quality Accurate and Correct (39)
Reliability and Consistency (13)
No Hallucination (4)
Balance between conciseness and thoroughness (14)
Answer Structure Complete with Examples (7)
Organised and Good Grammar (6)
Performance Efficiency (response time) (20)
Transparency (sources and confidence) (10)
Information Integrity Up-to-Date Information (4)
Data Security and Confidentiality (5)
Understanding Intent Understanding (10)
Domain and Contextual Understanding (2)
Usability Ease of Use (5)
Versatility Integration with Other Tools (1)
Adaptability to Contexts and Workflow (1)
Adaptability to Communication (2)
IV-E Frustration triggers in software engineering
We ask our participants to describe one or more situations of where they felt frustrated when interacting with an LLM. Regardless of the area of Software Engineering (industry or academia), we observed patterns of triggers that cause the LLM to fail to meet the software engineer’s expectations and hence cause frustration. In the following, we describe the triggers of a frustrating interaction between software engineers and LLMs.
Incorrect answers and hallucinations: One of the most common causes of frustration was continuously receiving incorrect or hallucinated responses from the LLM. While receiving such responses was referred to as “Annoying” and “Disappointing”, the frustration is caused when these responses are received continuously even when the engineers change their prompt or try to correct the LLM. Our participants described such situations as LLMs being stubborn and insisting on the an incorrect or hallucinated answer.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 3: “After several corrections, and repeating the prompt in different manners, it decided to reiterate the same wrong response.”
The definition of incorrectness varied among participants. Some examples of incorrect answers are uncompilable or buggy code, incorrect explanation of error message or incorrect factual information that were verified using other sources (e.g., an expert or a search engine). Hallucination was described as nonsense explanations, refernces that do not exist, functions and syntax in a code that do not exist.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 38: “During a coding problem, I was looking for the usage of a specific function in a library. I was frustrated when it provided a different [function] (which did not work or even exist).”
Intent not understood: Frustrations (or emotions that can evolve to frustration) are also caused when the software engineers get the impression that the LLM did not understand their prompt. Not understanding can be reflected on the response with an irrelevant response to the initial question.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 8: “I asked CodeLlama to explain the code of a function in a library with many lines, and it proceeded to explain a different function from a different library.”
However, a few of the participants pointed out that such misunderstandings can be caused by the poor phrasing or the lack of context in the prompt.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 5: “I provide short and maybe unclear prompts [then] I usually get irrelevant responses. The better the prompt, the better response.”
Personal preferences unmet: Many of our participants pointed out that they get frustrated with reasons that are related to their personal preferences. Some aspects of certain LLMs (for example, the answer structure) can be annoying to some engineers, and when the frequency of interactions with the LLM increases, the annoyance turns into frustration. For example, two of participants pointed out that the way that ChatGPT apologized every time they tried to correct the LLM or send a follow-up prompt was a source of frustration. Others were frustrated with how the LLM they use only provides answers that are long, or only in a bullet points. Forcing the LLM to structure an answer that aligns with their preferences required specifying many requirements and constraints in the prompt.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 35: “[I get frustrated] when ChatGPT gives to long answers. I quite often ask things that can be answered with a short sentence, but still I get half a page of ramblings back.”
Limitations of the LLM: Other frustration triggers were due to limitations of the LLM. Such limitations can lead the LLM to be unable to perform a specific task if it is not trained on to do so, or could be a limitation with the configurations of the LLM e.g., the size of the context window. Since software engineers cannot do anything to overcome these limitations, they get frustrated when they attempt to force the LLM to overcome limitations by prompting. For instance, a software developer can become frustrated when attempting to create a complex and large application using one prompt, but the LLM stops generating the response when the maximum output token is reached [^1] https://huggingface.co/docs/transformers/en/llm_tutorial resulting in an incomplete application code with missing lines of code with logic and syntax errors.
In these cases, the software engineer can use other LLMs that overcome these limitations rather than trying to prompt the same LLM differently.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 56: “[I got frustrated when] formatting of a table in latex, [had to] move to another LLM”
After describing their experiences, the participants indicated their perceived importance level of five aspects of an LLM, that is, Correctness, Hallucination, Understanding, Performance, and Ability to Answer, which can impact their overall experience. We report the importance levels that followed a 5-level Likert Scale in Figure 6. We found that the importance levels align with the frustration triggers we identified. Correctness, hallucination and understanding were assigned higher importance levels, indicating their high impact on the engineer’s experience when interacting with LLMs. Whereas the performance of the LLM (e.g., response time), and the ability to answer a question where deemed as less important.
Figure 6: Likert-scale results of the importance level of different aspects of the LLMs that can impact the user experience.
IV-F Unmet expectations’ impact on motivation
When the LLM fails to assist our participants, the participants’ motivation to complete their task was influenced mainly in three ways.
21.3% of our respondents reported that their motivation decreased when LLMs did not give the correct answer. Responses expressed frustration, stress, or disappointment, impacting participants’ willingness to continue. “My motivation decreases. I try to give it more context or constraints to perform the task”.
Some other participants commented that they eventually gave up. These respondents mentioned that after some effort, they decided to abandon the LLM and move on to other methods or stop altogether.
Another group was formed of 77% of the participants who were unfazed by the LLM’s failure, treating it as a non-critical tool and continuing with the task with their motivation not being affected.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 8: “My motivation is not affected, I just realise that the task will take longer”
In an interesting case, one participant mentioned that their motivation increased when the LLM did not assist them as intended, this was due to a desire to solve the problem manually or a learning experience.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 16: “I usually understand the problem a lot more so I want to complete the task more”
IV-G Suggestions and ways to reduce frustration
When asked what actions participants typically take after receiving an incorrect answer from an LLM, the majority (41 out of 62 participants) said they changed the prompt to try again. A smaller group (12 participants) reported providing feedback to the LLM, while a few (2 participants) said they did nothing. See Figure 7.
In the “Other” category, participants described various strategies. Some combine multiple actions, such as changing the prompt, switching to another LLM, or even reverting to traditional search methods like Google or Wikipedia. A few participants mentioned disengaging from the LLM entirely or writing the response themselves.
Figure 7: Actions done after receiving a wrong answer from LLMs
We asked our participant about the improvements they would like to see that would enhance their experience when using LLMs. They identified several areas; we grouped them and show the results in Table IV. The suggestions mainly align with the engineers’ expectation (See Table III). Regarding the quality of the outcome, the participants emphasized the importance of reducing hallucinations and improving response accuracy. They also commented on the need for enhanced analytical capacity to tackle more complex problems.
Regarding usability, participants expressed a preference for concise and short answers (unless they are prompted otherwise) and user-friendly interaction. They also stressed the importance of memory and learning capabilities, allowing LLMs to retain relevant information and build on past interactions. For understanding, prompt comprehension was considered essential, with users expecting LLMs to ask clarifying questions when necessary. Additionally, participants wanted transparent communication about the system’s limitations.
When it comes to the answer structure, users called for consistency in responses and better organisation of information. They also noted the need for LLMs to handle simple formatting tasks more efficiently. As for information quality, participants focused on the importance of correct references and proper citation of sources. They also mentioned concerns about confidentiality, urging transparency in how sources are used and shared. Finally, regarding versatility, users indicated the importance of handling more complex tasks and adapting to different user needs. Increased functionality to support a broader range of tasks was also crucial for improving overall user satisfaction.
TABLE IV: Improvements for Better User Experience
Themes Categories
Transparency (source, confidence) (16)
Information Integrity Trust and Data Security (4)
Relevant information (Up-to-date) (2)
State reasoning and assumptions (2)
Versatility Adaptability in communication (11)
Integration in project environment (2)
Extended Functionality (2)
Quality Improved Response Accuracy (6)
Reduced Hallucinations (4)
Less creative (3)
Understanding Context understanding (5)
Clarification questions (3)
Intent Understanding (2)
Answer Structure Elaborative answers on-demand (4)
Consistency in Responses (3)
Performance Higher memory utilization (4)
Efficiency (processing time) (1)
Human-related Training for engineers (2)
V Discussion
V-A Frustrations in the context of LLM interaction
We identified four main triggers that cause frustrations or feelings that can develop into frustration over the time. These causes are related to correctness and hallucinations, misunderstandings, personal preferences, and limitations of the used LLM. In code-related tasks, correctness and meeting personal preferences were important for the software engineers, since the absence of either can make the response from the LLM unusable. For instance, if the code is buggy or does not follow a standard practice within the software organization, it creates additional work. When the outcome is a piece of code and it is buggy, it requires the developer to perform debugging rather than implementation, which can be expensive in terms of time when the developer is working on code not written by them [fitzgerald2008debugging], especially, if it is written by an LLM which can produce less maintainable code [liu2024refining] and can be more frustrating than writing the code with no LLM assistance [weisz2022better]. In other cases, when the LLM is unable to provide a response due to a limitation of the model or a misunderstanding of the engineer’s intent (especially when the prompts is specific software-related research area), it can cause disruption to the workflow of the engineer when they attempt to make the LLM understand and provide the expected outcome, as we see in Figure 7, where most of the participants provide feedback, change the prompt, or change the LLM when they encounter an incorrect response. As a consequence, the progress of the software engineer can be delayed, making them feel as though they have spent longer than planned to reach their goal (e.g., completing a task), which enforces frustration and emotional strain [ceaparu2004determining, lazar2006workplace].
Furthermore, frustrating experiences among our participants resulted from several emotions other than frustration directly, such as anger and annoyance (See Figure 5). Such emotions discussed in studies that focused specifically on a GitHub Copilot [^2] https://github.com/features/copilot, an LLM-powered code assistant. Eshraghian at al. [eshraghian2024ai] shows that emotions such as frustration and anger can come from the feeling of a threat without being able to control it. For instance, practitioners where annoyed from the data privacy policy of GitHub Copilot, which has a threat of leaking confidential data (e.g., production code), with very little control over it (i.e., to use the LLM, they need to accept the policy). Frustrations caused by GitHub Copilot also centred around usability [usability2023frustration], where providing code suggestions, especially large ones, when they are not needed made developers frustrated. These Copilot-specific frustration causes were not observed specifically among our participants since very few of them used GitHub Copilot (See Figure 4). However, providing too large code suggestions and the need for the LLM to use the code data fit under our defined frustration triggers “failing to meet personal preferences” and “limitations of the LLM” respectively.
We distinguish the frustration triggers that we identified from frustrations that can be caused by technologies in other domains. In the general domain, frustrations caused by tools and technologies are due to performance issues (e.g., system crash) [lazar2006workplace], which we did not observe among our participants. While our participants mentioned that they expect the LLMs to be stable and efficient in terms of speed of response (See Table III), they were not frustrated because of any performance issues. We argue that this is due to recent and popular LLMs being more stable and efficient. Furthermore, technologies adopted in the medical domain are usually heavily relied on for clerical tasks (e.g., documentation), and frustrations can occur when the user interface is complex [opoku2015user], but are also associated with the nature of emotionally exhausting the work environment, which can enable frustration when working with technology at work [tawfik2021frustration]. We observed among our participants that LLMs are not essential for them to complete their tasks as they can rely on their expertise or use other alternatives (See Section IV-F). While in other domains, technologies can be more crucial and relied on e.g., using a technology to document a patient’s state.
Takeaway: The four main frustration triggers come from potential things that can require extra processing time (e.g., modifying an output) and can disrupt the workflow of the engineer.
V-B Impact of frustrating experiences on motivation
While most of our participants expressed frustration (or similar emotions), their motivation to complete the task was not necessarily affected. Our results showed that participants often felt demotivated when an LLM failed to meet their expectations, such as when it did not assist them as intended, frequently leading to frustration. However, most of those who felt frustrated reported that their motivation remained intact, likely due to perceiving the LLM as merely one tool among others in achieving their goal. When an LLM could not provide the necessary assistance, participants commonly resorted to alternative solutions, such as using search engines like Google or relying on their expertise to complete the task independently. These observations align with findings by Franca et al. [francca2014motivated], who explored the connection between motivation and the satisfaction and happiness of software engineers, revealing that happiness slightly overlaps with but does not correlate with motivation.
These findings suggest that, despite facing emotional challenges, software engineers maintain their motivation to persist with demanding tasks.
[…]
skipabove=2pt, hidealllines=true, leftline=true, innerlinewidth=2pt, innerlinecolor=quote, linewidth=0pt, backgroundcolor=quote!10 {mdframed}
Participant 39: “Sometimes the LLM hallucinates and puts me in loops. When I realize this, I resolve it myself using my human knowledge and 15 years of experience in the development field.”.
Although frustration may not directly impact motivation, it remains a critical factor to consider due to its known connection with burnout. Sustained frustration is still significant as it contributes to emotional strain [tawfik2021frustration] —a known precursor to burnout in high-demand professions like software engineering. Thus, understanding and managing frustration, even when the motivation appears unaffected, is essential in supporting the long-term well-being of software engineers.
Takeaway: Although frustration can be present in software engineers when LLMs fail to meet expectations, it generally does not diminish their motivation to complete tasks.
V-C Towards a less-frustrating user experience
Several studies focus on enhancing LLMs; however, it is essential to recognise that these systems will likely always have room for improvement. Therefore, our focus in this study shifts to the human element. We propose enabling users with the knowledge and skills to navigate LLMs effectively to mitigate frustration and improve overall user experience. The goal is to address frustration triggers that arise during their use. Specifically, triggers such as misunderstanding of the intent, unsatisfying personal preferences, or even getting incorrect answers can be minimised by prompt engineering the query before sending it to the LLM. Prompt engineering can involve using prompt techniques (such as Few-shot learning), including relevant contextual information (e.g., description of the system) or constraints about the output (e.g., the output structure). Other frustration triggers, such as hallucinations and limitations of LLMs, can be mitigated if the engineers know the LLM’s capabilities and limitations and align their expectations. Since unmet expectations are among the primary frustration triggers[levine2022unhappy] (see Section IV-D), providing software engineers with training on effective usage and clear understanding of LLM capabilities to set realistic expectations can help reduce disappointment and enhance user satisfaction.
Additionally, raising awareness about potential frustration triggers is important; engineers can manage their reactions accordingly if they recognise the likelihood of frustration in certain situations. For instance, using coping strategies rather than repeatedly attempting to elicit a perfect answer from the LLM. Therefore, to address the previous points, we suggest that software engineers (both in academia and industry) need training on how to use LLMs safely. This idea was also discussed by Barman et al. [barman2024beyond] where they propose providing guidelines for LLM users to know how to interact with different LLMs, for instance, if it is appropriate to generate artefacts or only to get some guidance. Our participants commented that they were very familiar with LLMs; however, most reported frustration, raising questions about whether they truly understand how to leverage LLMs effectively. Familiarity does not necessarily equate to proficiency[ladwig2012perceived], stressing the need for improved training and guidance on optimal usage strategies. A complete understanding of LLM capabilities and limitations can help users to manage their expectations.
Takeaway: Training on effective prompt engineering and a clear understanding of LLM capabilities and limitations to set realistic goals are important to address frustration triggers.
V-D Threats to Validity
In this section we explain how we addressed this study’s threats to validity.
Internal Validity: To ensure internal validity, we considered several biases and employed mitigations accordingly. Self-selection bias: Participation in the survey was voluntary; hence, individuals with particularly strong positive or negative experiences with LLMs might be overrepresented, skewing the data. To mitigate this, we tried to recruit diverse participants across different experience levels, regions, and fields. To address social desirability bias, we collected anonymous data by including a statement at the beginning of the survey and avoiding personal questions. This approach aimed to prevent participants from feeling pressured to align their responses with what they perceived as socially or professionally acceptable. This could lead to underreporting frustration to appear more competent with new technologies.
External Validity: To address external validity and avoid overrepresenting certain groups or regions, we targeted respondents from different countries and several fields within software engineering. Similarly, we employed stratified sampling to ensure a balanced representation across demographics, skill levels, and industries.
Construct Validity: We operationalised key concepts like frustration and hallucination to guarantee construct validity, adding their definitions. Additionally, we provided examples throughout the survey to clarify the scenarios we were exploring. Finally, we explained the Likert scale by adding information on how to measure each level. Further, we piloted the survey to ensure clarity. We used the feedback to fix ambiguous questions, clarify terms, and minimise confusion.
VI Conclusion
This study focused on the emotional strains, particularly frustration, experienced by software engineers when interacting with Large Language Models (LLMs) and not being assisted as intended. By identifying main triggers, such as the correctness and reliability of responses and issues related to personalisation, we emphasise that understanding the emotional impacts of LLM use in software engineering is important. This study’s insights bring attention to the potential risks to productivity and mental health if emotional responses go unaddressed. Future research should further explore the psychological implications of LLM use, focusing on sustainable strategies to support the well-being of software engineers and optimise their user experience with AI tools.